OpenStudy (anonymous):

Help. find the vertex and maximum value of f(x) f(x)=-2/7(x-4)^2+3

OpenStudy (raden):

the general form quadratic equation is f(x) = a(x-h)^ + k with (h,k) as the vertex and k as the max value

OpenStudy (raden):

so, from the equation we can look its answer so,, if given f(x)=-2/7(x-4)^2+3 it means the vertex = (4,3) and the max value = 3